# Plan2Align: Predictive Planning Based Test-Time Preference Alignment in Paragraph-Level Machine Translation

### v1

@@ B.1.2 RAIN
B.1.3 Vanilla MPC
Vanilla MPC applies MPC directly to paragraph-level machine translation (MT) by selecting the best output from the previous iteration and refining it. We conduct six iterations, with performance results presented in Table 6. The best performance for each language pair (iteration = 1, 2, 4 for zh→en, zh→ru, and zh→de, respectively) is reported in Table 1. The results show that even in the zh→en translation direction, performance continues to decline over iterations. For more challenging language pairs (zh→ru and zh→de), LLMs struggle to extract useful contexts, preventing sustained improvement and leading to unstable performance.
[…]
zh-en 1 89.79 58.40 17.30 4.55
2 88.00 56.26 16.45 6.42
3 83.92 53.08 15.11 10.70
4 84.66 53.60 15.26 9.93
5 84.93 54.03 15.48 9.68
6 84.48 52.91 15.12 10.08
zh-en 1 73.80 30.83 5.54 20.00
2 82.83 35.18 7.19 10.26
3 77.08 34.08 6.15 16.40
4 79.56 33.25 6.21 13.74
5 78.35 33.31 6.34 15.35
6 78.35 33.37 5.43 15.24
zh-en 1 83.06 39.00 10.89 11.19
2 80.88 36.74 8.95 13.03
3 79.79 36.72 9.70 14.51
4 83.67 38.46 10.82 10.63
5 82.72 38.54 10.25 11.40
6 79.35 35.87 9.54 14.87
Table 6: Performance of Vanilla MPC across different iterations.
B.2 Training-Time Alignment Methods
We choose LLaMA3.1-8B-Instruct as our experiments’ backbone model, including the reward model and training-time methods. We utilize one NVIDIA RTX 6000 Ada Generation GPU to train models with the LLaMA Factory library [^3] https://github.com/hiyouga/LLaMA-Factory (Zheng et al., 2024). For the training setups, the SFT model is trained on preferred data from the preference dataset, while the Reward Model, DPO, and SimPO are trained on the full preference data that includes inputs, good translations, and bad translations. All models are configured with an identical set of hyperparameters: they utilize the AdamW(Loshchilov and Hutter, 2019) optimizer and feature gradient accumulation steps fixed at 8, with a cutoff length of 2048. The maximum allowable gradient norm is restricted to 1.0 to ensure training stability. All training procedures are conducted using bf16 precision, each model undergoes a single epoch of training with a batch size of 2, and the LoRA(Hu et al., 2022) rank and LoRA alpha parameters are set to 16 across all models.
The learning rates for the models are: Rewarg Model at 1e-5, SFT at 2e-5, and both DPO and SimPO at 1e-6 as suggested by (Rafailov et al., 2023) and (Meng et al., 2024). Additionally, SimPO utilizes a beta and gamma of 2.5, which aligns with its setup, whereas SFT and DPO use a beta of 0.1 also followed by (Rafailov et al., 2023). When Plan2Align serves as a generic method, we also plan to connect the method for multi-agent (i.e., cascaded) based translation system Hu et al. (2024) in the future.
Appendix C Implementation Details of Plan2Align
@@
In this section, we describe our parameter settings and prompt design, as well as the impact of different prompts on the final translation generation. Our full approach is outlined in Algorithm 1.
Algorithm 1 Plan2Align
@@
0:  Source $x$, number of MPC iterations $N$, base model $\mathrm{LM}$, reward model $r$
0:  Translation result $y_{final}$
1:  Initial buffer $\text{Initial buffer }B$
2:  for $k=0\to N-2$ do
[…]
5:     for all $(w_{x},w_{y})\in(W_{x},W_{y})$ do
6:        $\mathrm{reward}\leftarrow r(w_{x},w_{y})$ // Compute rewards for each window pair
7:        Update the best window at each position to put into Buffer $B$ // Window selection
8:     end for
[…]
10:  end for
11:  return  $y_{final}$
C.1 Parameter Setting
[…]
C.2 Prompt Design
@@
Here is the prompt configuration for generating translations in Plan2Align: Figure 4 displays our prompt template for the self-rewriting task within Plan2Align. We employ three distinct system prompts that yield three types of translations: sentence-by-sentence translation, precise translation, and imaginative translation. These prompts cater to different translation needs within the context buffer, allowing us to extract the most effective parts. Additionally, we use a context prompt to ensure the quality of the translations produced. Figures 5 and 6 showcase the prompts used for generating the final translations in Plan2Align. The distinctions between these two prompts, Reference and Annotation, are further explored in Section C.2.1.
@@
Figure 4: Prompt template used for self-rewriting task in Plan2Align and an actual example.
@@
Figure 5: Prompt used for generating final translation in Plan2Align (the Reference version of the prompt in Section C.2.1) and an actual example.
Figure 6: Applying the concept of the self-rewriting task to generate the final translation prompt (the Annotation version of the prompt in Section C.2.1).
C.2.1 Impact of Final Translation Task
In Plan2Align, final translation generation differs from context buffer updates. The final translation is produced by assembling buffer-stored contexts in source sentence order, forming a reference translation (Reference). Conversely, during context buffer updates, translated sentences are annotated beneath their corresponding source sentences (Annotation). Table 7 shows that the Reference approach outperforms Annotation in zh→ru and zh→de. However, for zh→en, the Annotation method generally performs better. This difference arises because annotating each source sentence with context buffer information and refining segments independently, although not requiring explicit alignment, demands higher semantic understanding. This makes it effective for zh→en but burdensome for the other two translation directions. In contrast, the Reference approach offers greater flexibility in refining the entire paragraph.
[…]
zh-en Annotation 94.02 67.83 22.16 0.38
Reference 94.13 67.06 21.17 0.20
zh-ru Annotation 80.58 35.42 10.09 12.83
Reference 91.23 40.11 10.69 1.49
zh-de Annotation 80.25 37.27 9.09 13.54
Reference 92.60 44.62 10.69 1.44
Table 7: Performance of Plan2Align across two different final translation tasks.
Appendix D Evaluation Sanity Check
@@
Appendix E MetricX Scoring Failure Cases
Here are three failure cases for metricX-24, each illustrated with different models and language pairs:

### v2

@@ F.1.2 RAIN
F.2 Training-Time Alignment Methods and Reward Model
We adopt LLaMA3.1-8B-Instruct as the backbone model for all experiments, including the reward model and training-time alignment methods. Model training is conducted using a single NVIDIA RTX 6000 Ada Generation GPU, with the implementation based on the LLaMA Factory library [^6] https://github.com/hiyouga/LLaMA-Factory (44).
For training setups, the SFT model is trained solely on the preferred responses from the preference dataset, while the Reward Model, DPO, and SimPO are trained on the full preference data, which includes the input, chosen, and rejected responses, as detailed in Appendix C.
All models are trained using identical hyperparameters: the AdamW optimizer (20) with gradient accumulation steps set to 8, a sequence cutoff length of 2048 tokens, and a maximum gradient norm clipped at 1.0 for stability. Training is performed using bf16 precision, with a batch size of 2, for one epoch. We apply LoRA (12) with a rank of 16 and alpha of 128 across all models.
The learning rates are configured as follows: 2e-5 for the Reward Model and SFT, and 5e-6 for both DPO and SimPO, in accordance with recommendations from Rafailov et al. (27) and Meng et al. (22). Additionally, SimPO is trained with a gamma value of 0.4, while DPO uses a beta value of 0.1, consistent with the original settings proposed in their respective works.
Appendix G Implementation Details of Plan2Align and Vanilla MPC
@@
In this section, we describe our parameter settings and prompt design. Our full approach is outlined in Algorithm 1.
Algorithm 1 Plan2Align: Test-Time Preference Alignment via Predictive Planning
@@
0:  Input prompt $x$, base LLM $\mathcal{M}$, reward model $\mathcal{R}$, number of iterations $T$, rollouts per iteration $K$
0:  Final response $y^{\text{final}}$
1:  Initialize reference buffer $\mathcal{B}\leftarrow\emptyset$
2:  Initialize candidate response set $\mathcal{Y}\leftarrow\emptyset$
3:  Generate initial response $y^{(0)}\leftarrow\mathcal{M}(x)$
4:  Add $(y^{(0)},\mathcal{R}(y^{(0)}))$ to $\mathcal{Y}$
5:  Add segments from $y^{(0)}$ to buffer $\mathcal{B}$
6:  for $t=1$ to $T$ do
7:     Sample $K$ candidate responses $\{y_{k}^{(t)}\}_{k=1}^{K}\leftarrow\mathcal{M}(x,\mathcal{B})$Rollouts}
8:     for each $y_{k}^{(t)}$ do
9:        Compute reward $r_{k}\leftarrow\mathcal{R}(y_{k}^{(t)})$
10:        if task is not translation then
11:           Add $(y_{k}^{(t)},r_{k})$ to $\mathcal{Y}$
12:        end if
13:        Decompose $y_{k}^{(t)}$ into segments $\{s_{j}\}$
14:        for each segment $s_{j}$ do
15:           Compute segment reward $r_{j}\leftarrow\mathcal{R}(s_{j})$
16:           if $r_{j}>$ threshold then
17:              Add $s_{j}$ to buffer $\mathcal{B}$ at position $j$
18:           end if
19:        end for
20:     end for
21:     if task is translation then
22:        Construct final rewriting prompt using all segments in $\mathcal{B}$
23:        Generate final translation $y^{\text{trans}}\leftarrow\mathcal{M}(x,\mathcal{B}_{\text{all}})$
24:        Compute reward $r^{\text{trans}}\leftarrow\mathcal{R}(y^{\text{trans}})$
25:        Add $(y^{\text{trans}},r^{\text{trans}})$ to $\mathcal{Y}$
26:     end if
27:  end for
28:  Select $y^{\text{final}}\leftarrow\arg\max_{(y,r)\in\mathcal{Y}}r$
29:  return  $y^{\text{final}}$
G.1 Long-Form Responses Generation
G.1.1 Parameter Setting
To generate diverse translations for comparison, we design three distinct system prompts for the rewriting task, resulting in three translation outputs per iteration. The segment size is fixed at 3 sentences, and the number of iterations is set to 3. The reference buffer threshold is fixed at 3.
G.1.2 Rewriting Prompt Design
We present the prompt configuration used for generating translations in Plan2Align. As shown in Figure 8, our prompt template is designed for the rewriting task within Plan2Align. We utilize three distinct system prompts, each aimed at a specific type of improvement: making the text more concise and clear, more detailed and informative, or more balanced and well-structured.
G.2 Paragraph-Level Machine Translation
G.2.1 Parameter Setting
To generate diverse translations for comparison, we design three distinct system prompts for the rewriting task, resulting in 3 translation outputs per iteration. The segment size is fixed at 3 sentences, and the number of iterations is set to 3. The reference buffer threshold is fixed at 1, as the translation task already introduces randomness through sampling at different segment positions. More specifically, when constructing the rewriting input from the reference buffer, we adopt a sampling strategy to encourage diversity and uncover more valuable segments. As the number of iterations increases, the reference buffer accumulates more content. To leverage this growing pool, we adaptively increase the number of reference segments used by Plan2Align. In particular, we define the number of sampled segments as a function of the current iteration using the formula: sampled segment number = current iteration number + 5.
G.2.2 Rewriting Prompt Design
@@
We present the prompt configuration for generating translations in Plan2Align: Figure 6 displays our prompt template for the rewriting task within Plan2Align. We employ three distinct system prompts that yield three types of translations: sentence-by-sentence translation, precise translation, and imaginative translation. These prompts cater to different translation needs within the reference buffer, allowing us to extract the most effective parts. Figures 7 showcase the prompts used for generating the output translations in Plan2Align.
G.3 Implementation of Vanilla MPC
Vanilla MPC serves as a simplified ablation of Plan2Align, aiming to test whether repeated refinement alone—without segment-level planning—is sufficient to improve generation quality. At each iteration, the model performs multiple rollouts and selects the response or translation with the highest reward score, as determined by the reward model. This top-performing output is then passed to the language model as input for further refinement. After several iterations, the final output is selected from all responses based on reward scores.
We present the full algorithmic steps of Vanilla MPC in Appendix Algorithm 2.
System Prompts for Response Generation.
To encourage meaningful revisions in long-form response generation, we define three distinct system prompts. For fairness, they are closely aligned with those used in Plan2Align. As shown in Figure 8, these prompts are designed to guide the model toward generating improved outputs with varying qualities.
  • Improve the given response. Make it more concise and clear. Respond only with the improved answer.
  • Improve the given response. Make it more detailed and informative. Respond only with the improved answer.
  • Improve the given response. Make it balanced and well-structured. Respond only with the improved answer.
System Prompts for Paragraph-Level Translation.
To guide stylistic refinement in paragraph-level translation, we define three distinct system prompts, each corresponding to a different translation intent. As shown in Figure 6, these prompts encourage the model to revise previous translations with specific goals in mind.
  • You are a meticulous translator. Provide a literal, word-for-word translation that preserves the structure and meaning of each individual word.
  • You are a professional translator. Deliver a clear, formal, and precise translation that faithfully conveys the original meaning.
  • You are a creative and expressive translator. Render the text in a vivid and imaginative way, as if narrating a captivating story.
In addition, we provide a context prompt that presents the previously translated result:
  • In the previous evaluation, here is the result you have translated before: history_mt. Please consider this result and improve it to give a better translation.
This setup encourages the model to generate improved translations by leveraging prior outputs, guided by both explicit intent and contextual information.
Algorithm 2 Vanilla MPC for Test-Time Refinement
0:  Input $x$, base LLM $\mathcal{M}$, reward model $\mathcal{R}$, number of iterations $T$, rollouts per iteration $K$
0:  Final response $y^{\text{final}}$
1:  Initialize candidate set $\mathcal{Y}\leftarrow\emptyset$
2:  Generate initial response $y^{(0)}\leftarrow\mathcal{M}(x)$
3:  Add $(y^{(0)},\mathcal{R}(y^{(0)}))$ to $\mathcal{Y}$
4:  for $t=1$ to $T$ do
5:     Generate $K$ rollouts $\{y_{k}^{(t)}\}_{k=1}^{K}\leftarrow\mathcal{M}(y^{(t-1)},\text{prompt}_{t})$
6:     for each $y_{k}^{(t)}$ do
7:        Compute reward $r_{k}\leftarrow\mathcal{R}(y_{k}^{(t)})$
8:        Add $(y_{k}^{(t)},r_{k})$ to $\mathcal{Y}$
9:     end for
10:     Set $y^{(t)}\leftarrow\arg\max_{y}\mathcal{R}(y)$ over current $K$ rollouts
11:  end for
12:  return $y^{\text{final}}\leftarrow\arg\max_{(y,r)\in\mathcal{Y}}r$
@@ G.3 Implementation of Vanilla MPC
Figure 6: The template of rewriting prompt used for translation task in Plan2Align and an actual example.
@@
Figure 7: Prompt used for generating final translation in Plan2Align and an actual example.
Figure 8: The template of rewriting prompt used for response generation task in Plan2Align and an actual example.
Appendix H Evaluation Sanity Check
@@
H.1 Limitations of MetricX in Detecting Translation Errors
We conduct a preliminary study to assess whether state-of-the-art MT metrics can reliably detect under- and over-translation errors when applied to concatenated sentence-level outputs. Using MetricX-24 and its reference-free variant MetricX-24-QE from the WMT 2024 Metrics Shared Task (8), we evaluate three language pairs: en-de, en-es, and ja-zh. To stay within the 1024-token input limit of MetricX-24, we filter out overly long examples.
We simulate under-translation by removing sentences from the target side, and over-translation by removing sentences from both source and reference. We limit ourselves to removing at most two sentences since removing more would leave too few documents for reliable Kendall’s $\tau$ correlation calculations. We measure the performance of the metrics by computing Kendall’s $\tau$ correlations between the metric outputs and human evaluation scores. For cases with introduced under-/over- translate, we assign $25$ as the human-annotated MQM score for each , which is then converted into z-score in accordance with each human annotator’s scoring distribution. Finally we average the segment-level human judgment scores as the document-level scores.
As shown in Figure 9, both metrics exhibit limited sensitivity to these common error types. These findings motivate our use of CW-COMET and CW-KIWI, which explicitly account for cross-sentence coherence and alignment in long-form translation evaluation.
Figure 9: The Kendall’s $\tau$ correlations of MetricX-24 and MetricX-24-QE show limited sensitivity to over- and under-translation; sentences with more than three drops are insufficient to estimate correlations reliably.
H.2 MetricX Scoring Failure Cases
In addition to its limited ability to detect translation errors, we also find that MetricX-24 occasionally produces nonsensical judgments. Below, we present three failure cases from different models and language pairs to illustrate this issue.